# Calibrated Self-Rewarding Vision Language Models

### v3

@@ 2 Preliminaries
where $\pi_{\text{ref}}(y|x)$ represents the reference policy, i.e., language model after supervised fine-tuning.
@@ 4.2 Results
CSR Continuously Improves Model Performance over Iterations. In Figure 3, we report the average performance of LLaVA-1.5 7B and 13B models concerning the number of training iterations on comprehensive benchmarks, general VQA tasks, and hallucination benchmarks. In the experiment, the 7B model achieved an improvement of approximately 7.62% across all benchmarks through online iterative updates, while the 13B model saw an improvement of approximately 5.25%. According to the full results in Table 6 and Table 7 of Appendix A.5, the improvement is particularly significant on the LLaVAW and CHAIR benchmarks, with improvements of 8.9% and 49.50%, respectively. The results indicate that CSR is capable of incrementally improving model performance over iterations, demonstrating its effectiveness in self-improving the quality of generated preference data and leading to stronger modality alignment. The degree of improvement gradually becomes smaller, which is not surprising, indicating that the model is gradually converging.
@@
Base 65.96 67.56
Only RT 67.66 68.70
Only RI 66.77 68.23
CSR (Ours) 70.99 71.10
@@ 4.3 Analysis
Ablation Study. To validate the effectiveness of using the image-response relevance score ($R_{I}$) to complement the self-generated instruction following score ($R_{T}$), we specifically compare CSR with three variants: (1) without applying CSR on LLaVA 1.5 (Base); (2) using CSR with only the self-generated instruction following score (Only $R_{T}$); and (3) using CSR with only the image-response relevance score (Only $R_{I}$). The results are reported in Table 2. We first observe that CSR improves performance by jointly considering both the self-generated instruction following and image-response relevance scores. This verifies its effectiveness in enhancing modality alignment by calibrating the language-driven self-rewarding paradigm with visual constraints.
@@
Figure 5: The change in image relevance scores before and after employing CSR.
@@ 6 Related Work
Large Visual-Language Model Hallucination. Recently, the rapid development of visual-language alignment methods [19, 44, 45, 46] and LLMs [47, 48, 49, 50, 51] has significantly accelerated the progress of LVLMs, which extend LLMs with visual modalities and demonstrate impressive visual understanding by unifying the encoding of visual and text tokens [52, 53, 54, 55, 56, 57]. However, LVLMs still face the problem of hallucination [58, 59, 60], where generated text descriptions contradict the visual modality information. Various approaches have been proposed to address hallucination in LVLMs, including enhancing dataset quality for fine-tuning [61, 8, 62, 9], manipulating the decoding process [37, 63, 64, 65, 66, 67], and leveraging external closed-source models to facilitate post-hoc mitigation of hallucination [58, 68]. Though these approaches alleviate hallucination to some extent, they do not focus directly on improving modality alignment.
@@
Self-Improvement in Large Language Models. Self-improvement emerges as a powerful paradigm for LLMs to enhance themselves without significant external intervention. For example, self-rewarding and online alignment [71] propose a method that selects consistent answers generated by the model to fine-tune itself, thereby improving its reasoning ability. Similarly, [12] utilizes self-play to enhance the model’s performance by distinguishing its self-generated responses from those in human-annotated training data. Unlike prior methods that primarily target LLMs, CSR addresses the modality misalignment issue in LVLMs during the preference modeling process by introducing visual constraints, making it particularly well-suited for LVLMs.
@@ A.1.1 Hyperparameter Settings
Sentence-Level Beam Search. We configure our parameters as follows to ensure both diversity and quality in the sampled data. The num_beams_parameter, set to 5, determines the capacity of input at each search layer. Additionally, num_token_beams, also set to 5, ensures that each beam search returns 5 token-level search results. The eos_token_id is set to the token for a period, effectively controlling the sentence-by-sentence generation process. The max_length parameter, set to 1024, prevents truncation errors and infinite repetitions by controlling the maximum length, while max_new_tokens, set to 74, limits the maximum length of newly generated content to avoid exceeding the CLIP encoding limit.
@@ A.5 Additional Experiments
In this subsection, we provide a comparison of CSR with other state-of-the-art models, a performance comparison of different CSR iterations, a comparison of hallucinations in different CSR iterations, and validation experiments of CSR on other models. Experiments strongly demonstrate the effectiveness of CSR.
@@ Appendix B Proofs
We complete the proof. ∎

### v4

@@ 2 Preliminaries
where $\pi_{\text{ref}}(y|x)$ represents the reference policy, i.e., the language model after performing supervised fine-tuning.
@@ 4.2 Results
CSR Continuously Improves Model Performance over Iterations. In Figure 3, we report the average performance of LLaVA-1.5 7B and 13B models concerning the number of training iterations on comprehensive benchmarks, general VQA tasks, and hallucination benchmarks. To facilitate score calculation, we first calculated an average score on a 100-point scale by adjusting the original values: MMEP was divided by 16, and MMEC was divided by 4, corresponding to the number of categories in MME. Additionally, since a lower CHAIR value indicates better performance, we standardized all metrics to follow a higher is better approach by transforming the CHAIRS and CHAIRI metrics into 100 - CHAIRS and 100 - CHAIRI. We then calculated the average score by averaging these standardized values, which were used to compute the average percentage increase. In the experiment, the 7B model achieved an improvement of approximately 7.62% across all benchmarks through online iterative updates, while the 13B model saw an improvement of approximately 5.25%. According to the full results in Table 6 and Table 7 of Appendix A.5, the improvement is particularly significant on the LLaVAW and CHAIR benchmarks, with improvements of 8.9% and 49.50%, respectively. The results indicate that CSR is capable of incrementally improving model performance over iterations, demonstrating its effectiveness in self-improving the quality of generated preference data and leading to stronger modality alignment. The degree of improvement gradually becomes smaller, which is not surprising, indicating that the model is gradually converging.
@@
Base 66.61 68.08
Only RT 68.46 68.12
Only RI 67.49 69.23
CSR (Ours) 72.39 71.95
@@ 4.3 Analysis
Ablation Study. To validate the effectiveness of using the image-response relevance score ($R_{I}$) to complement the self-generated instruction following score ($R_{T}$), we specifically compare CSR with three variants: (1) without applying CSR on LLaVA 1.5 (Base); (2) using CSR with only the self-generated instruction following score (Only $R_{T}$); and (3) using CSR with only the image-response relevance score (Only $R_{I}$). The results are reported in Table 2. We first observe that CSR improves performance by jointly considering both the self-generated instruction following and image-response relevance scores. This verifies its effectiveness in enhancing modality alignment by calibrating the language-driven self-rewarding paradigm with visual constraints. Additionally, we further conduct the analysis on the change of $\lambda$ in Eqn (4) and found that incorporating external visual scores to calibrate the model’s rewarding process effectively enhances performance (see detailed results in Appendix A.5.)
@@
Figure 5: Image relevance scores before and after employing CSR.
@@
4.4 Case Study
In this section, we use LLaVA-1.5 13B as an example to illustrate changes in the model’s own responses during CSR iterations and the preference data sampled in the CSR learning process, with hallucinations and errors highlighted in red. The results are shown in Figures 7 and 9, respectively. As shown in Figures 7, with each iteration of CSR, hallucinations in the model’s responses noticeably decrease. This indicates that CSR effectively refines the model’s preferences through iterative preference learning, making the model’s responses more accurate. In Figure 9 of the Appendix, we present a sampled preference data pair from the model during the CSR learning process. It can be seen that through CSR, the model not only gradually refines its own preferences but also obtains high-quality preference data pairs without human annotation.
Figure 7: Two cases selected from the CSR-generated datasets.
@@ 6 Related Work
Large Visual-Language Model Hallucination. Recently, the rapid development of visual-language alignment methods [19, 44, 45, 46, 47, 48, 49] and LLMs [50, 51, 52, 53, 54] has significantly accelerated the progress of LVLMs, which extend LLMs with visual modalities and demonstrate impressive visual understanding by unifying the encoding of visual and text tokens [34, 55, 56, 57]. However, LVLMs still face the problem of hallucination [58, 59], where generated text descriptions contradict the visual modality information. Various approaches have been proposed to address hallucination in LVLMs, including enhancing dataset quality for fine-tuning [60, 8, 61, 9], manipulating the decoding process [37, 62, 63, 64, 65, 66], and leveraging external closed-source models to facilitate post-hoc mitigation of hallucination [58, 67, 68, 69, 70]. Though these approaches alleviate hallucination to some extent, they do not focus directly on improving modality alignment.
@@
Self-Improvement in Large Language Models. Self-improvement emerges as a powerful paradigm for LLMs to enhance themselves without significant external intervention. For example, self-rewarding and online alignment propose a method that selects consistent answers generated by the model to fine-tune itself [73, 74] , thereby improving its reasoning ability. Similarly, Chen et al. [12] utilizes self-play to enhance the model’s performance by distinguishing its self-generated responses from those in human-annotated training data. Unlike prior methods that primarily target LLMs, CSR addresses the modality misalignment issue in LVLMs during the preference modeling process by introducing visual constraints, making it particularly well-suited for LVLMs.
@@ A.1.1 Hyperparameter Settings
Sentence-Level Beam Search. We configure our parameters as follows to ensure both diversity and quality in the sampled data. The num_beamsparameter, set to 5, determines the capacity of input at each search layer. Additionally, num_token_beams, also set to 5, ensures that each beam search returns 5 token-level search results. The eos_token_id is set to the token for a period, effectively controlling the sentence-by-sentence generation process. The max_length parameter, set to 1024, prevents truncation errors and infinite repetitions by controlling the maximum length, while max_new_tokens, set to 74, limits the maximum length of newly generated content to avoid exceeding the CLIP encoding limit.
@@ A.5 Additional Experiments
In this subsection, we provide a comparison of CSR with other state-of-the-art models, a performance comparison of different CSR iterations, a comparison of hallucinations in different CSR iterations, validation experiments of CSR on other models, ablation study on $\lambda$ in Eqn (4), and the relationship between reward score and average performance score. Experiments strongly demonstrate the effectiveness of CSR.
For the ablation study on $\lambda$ in Eqn (4), our training settings are consistent with Table 1, with three rounds of iteration. The experimental results in Table 9 show that as the value of $\lambda$ increases, the model’s performance on various benchmarks improves. This suggests that calibrating the model’s rewarding process using the visual score can enhance the preference learning process, thereby boosting performance.
@@
Table 9: Performance comparison of CSR on LLaVA-1.5 7B with different $\lambda$ values on various benchmarks.
($\lambda=0.1$) 1508.6 369.3 60.0 66.7 64.9 31.6 70.0 54.0 62.0 86.90 40.8 10.2
($\lambda=0.5$) 1515.4 364.5 60.1 68.2 64.9 32.4 69.7 54.0 62.1 86.90 28.2 6.7
($\lambda=0.9$) 1524.2 367.9 60.3 71.1 65.4 33.9 70.7 54.1 62.3 87.01 21.0 6.0
Table 10: Reward score and average performance score across multiple iterations of CSR on LLaVA-1.5 7B.
Iteration Iter-1 Iter-2 Iter-3 Iter-4 Iter-5
Chosen reward 0.4885 0.5040 0.5052 0.5055 0.5066
Rejected reward 0.4551 0.4788 0.4789 0.4794 0.4799
Avg performance score 66.61 71.02 71.74 72.09 72.24
@@ Appendix B Proofs
We complete the proof.
Figure 9: A case including both self-generated preferred and dispreferred responses.
Appendix C Limitations
Due to limitations in computing resources, we conducted only three iterations of CSR. Additionally, our experiments were confined to 7B and 13B models. This restriction prevents us from determiing if our method adheres to a scaling law. We hope to continue iterative training in the future and to train larger models, given access to more computing resources, to explore the upper limits of our method.
Appendix D Broader Impacts
Our approach requires no additional human annotations and significantly enhances model performance using the model itself. Technically, our method may inspire more researchers to explore how multimodal models can learn from themselves. From a societal impact perspective, our method significantly reduces hallucinations in LVLMs, a major factor affecting the application of AI in real-world scenarios. Our approach promotes more responsible use of LVLMs. However, it is important to note that while our method significantly reduces hallucinations, they still occur. Therefore, it is crucial to employ various measures to ensure safety and stability when applying this approach in real-world scenarios.